the difference between two numbers is 11. Twice the smaller plus three times larger is 123. What are the numbers?
please help set up

Answer by Nate(3500)

l - s = 11
and
2s + 3l = 123
combine:
2l - 2s = 22
+
2s + 3l = 123
=
5l = 145
l = 29
Remember: l - s = 11 : 29 - s = 11 : -s = -18 : s = 18
l = 29
s = 18